A dipole with a dipole moment of strength p= 5.64*10^-4 Cm is located D= 0.71 m away from a grounded plane conducting surface. The direction of the dipole moment is perpendicular to the conducting surface. Calculate the strength of the translational force on the dipole with positive values indicating a force away from the surface and negative values a force toward the surface. The displace between the charges forming the dipole is much smaller than the distance of the dipole from the surface.
1
Expert's answer
2019-04-22T11:12:56-0400

The translational force F is given by formula


"F=E \times D (1)"

where E is the strength of the field


The strength of the field


"E=\frac {1} {4\pi ɛ_0} \frac {p} {D^3} \sqrt{1+3 \cos^2 \theta} (2)"

where ɛ0 is the electric constant

In our case,


"\theta=90°"

We put (2) in (1):


"F=\frac {1} {4\pi ɛ_0} \frac {p} {D^2} \sqrt{1+3 \cos^2 \theta} (3)"

Using (3) we got:


"F=4 \times 10^7 N"

Answer:

"4 \times 10^7 N"